OpenStudy (anonymous):

multiply (x-11)(x-2i)(x+2i)?

OpenStudy (anonymous):

Foil method!!

OpenStudy (anonymous):

Or how I like to call it. The rainbow method.

OpenStudy (anonymous):

first start with \[(x-2i)(x+2i)=x^2-(2i)^2=x^2-4i^2=x^2-4\times (-1)=x^2+4\] wow i really dragged that one out

OpenStudy (anonymous):

oh okay, I forgot about that method.

OpenStudy (anonymous):

rainbow good foil bad

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

what?

OpenStudy (apoorvk):

lol

OpenStudy (anonymous):

I took your seriously for a second there.

OpenStudy (anonymous):

http://www.youtube.com/watch?v=PW_97-R3S4M

OpenStudy (anonymous):

well ty for helping